OpenStudy (cedric):

Choose the correct simplification of the expression (4a4b)2. 16a8b 16a8b2 16a6b2 8a8b2

OpenStudy (anonymous):

what r u saying

OpenStudy (anonymous):

\((4a^4b)^2\) is my guess

OpenStudy (anonymous):

if so, the square everything, you get \(4^2\times (a^4)^2\times b^2=16a^8b^2\)